# On the resolution of categorical symmetries in (Non-) Unitary Rational CFTs

### v1

@@ Abstract
We explore several aspects of the categorical symmetry-resolved entanglement entropy (SREE) in two-dimensional Rational Conformal Field Theories (RCFTs) and express it directly in terms of the modular data of the theory. Motivated by the SymTFT construction of Choi:2024wfm, we provide a general formula for SREE that applies to symmetric (weakly/strongly) and cloaking boundary conditions as well as for fusion rings with multiplicities. We check the formula against several explicit examples. Additionally, we study symmetry resolution for both categorical and invertible symmetries in (non-)diagonal RCFTs and comment on the subtleties that arise in these cases. Finally, we extend our analysis to diagonal non-unitary RCFTs, focusing on theories with generalized Haagerup–Izumi modular data, and find full agreement with the given formula.
@@ 1 Introduction
Symmetries are omnipresent in nature. In physics, symmetry plays a crucial role in understanding the dynamics of quantum field theory and provides a framework for going beyond. According to Wigner Wigner:1959, each symmetry is associated with a unitary operation that has an inverse, i.e., group-like. Gaiotto et al. Gaiotto:2014kfa showed that the generalized global symmetry of a theory is associated with a topological defect operator, and there are plenty of examples found in condensed matter systems that realize this generalized global symmetry McGreevy:2022oyu. One of the classes of these generalized symmetries is famously known as non-invertible symmetry Bhardwaj:2017xup; Chang:2018iay; Thorngren:2019iar; Heidenreich:2021xpr; Thorngren:2021yso; Choi:2022jqy; GarciaEtxebarria:2022vzq; Cordova:2022ieu; Cordova:2022ruw; Choi:2022fgx; Choi:2022rfe; Apruzzi:2022rei; Shao:2023gho; Seiberg:2023cdc; Schafer-Nameki:2023jdn; Bhardwaj:2023wzd; Bhardwaj:2023ayw; Bhardwaj:2023fca; Choi:2023pdp; Lin:2023uvm; Benedetti:2023owa; Iqbal:2024pee; Choi:2024rjm; Seiberg:2024gek; Bhardwaj:2024wlr; Bhardwaj:2024kvy; Putrov:2024uor; Shao:2025mfj; Shao:2025qvf  [^1] The references mentioned here is by no means exhaustive. Interested readers are referred to some of these reviews and lecture notes and their citations McGreevy:2022oyu; Shao:2023gho; Schafer-Nameki:2023jdn for more references., which can not be described by a unitary operation, more specifically, it does not have an inverse. In mathematics literature, it is known as a category Bartsch:2023wvv; Costa:2024wks. For these kinds of categorical (non-invertible) symmetries, associated topological operators $\{\mathcal{L}_{a}\}$ obey fusion rules of a tensor category rather than a group law, as shown below,
@@
There are numerous examples that demonstrate this kind of categorical symmetry exists in various branches of physics, including high-energy theory, mathematical physics, and condensed matter physics, particularly in low-dimensional quantum systems that are useful for quantum computation Preskill:1999he; Kitaev:2005dm; Kitaev:1997wr; PhysRevLett.96.110405; Aharonov:2005nkd; Feiguin:2006ydp; Dennis:2001nw; McGreevy:2022oyu; PRXQuantum.5.037001; Okada:2024qmk; Li:2024gwx; Bhardwaj:2024ydc; PhysRevX.15.011058; Li:2025bgo; KNBalasubramanian:2025vum. Before delving into the main discussion of the paper regarding the categorical symmetry resolution of the entanglement entropy, we briefly discuss the representation of categories and the importance of dealing with the Drinfeld centre or the quantum double $Z(\mathcal{C})$ of a given category $\mathcal{C}\,.$
@@
Parallel to recent advances in the characterization of (non-)invertible symmetries, a complementary line of research has focused on the role of symmetries in shaping entanglement structures in quantum field theories (QFTs). Entanglement serves as a fundamental measure of quantum correlations and phases of matter, and provides a particularly sensitive probe of underlying symmetry principles, including generalized and non-invertible symmetries. In this context, symmetry-resolved entanglement entropy (SREE) Goldstein:2017bua; german3; Belin:2013uta; Islam:2015mom; doi:10.1126/science.aau0818; FG; Murciano:2019wdl; Murciano:2020lqq; Parez:2021pgq; Murciano:2020vgh; Zhao:2020qmn; ourPartI; Horvath:2021rjd; Ares:2022gjb; DiGiulio:2022jjd; Gaur:2023yru; Kusuki:2023bsp; Northe:2023khz; Gaur:2024vdh; Banerjee:2024ldl [^2] For more references and details, interested readers are referred to the following review of symmetry-resolved entanglement measure Castro-Alvaredo:2024azg. has emerged as a refined diagnostic that decomposes the total entanglement into contributions associated with individual symmetry sectors, thereby elucidating how symmetry structures govern the distribution of quantum information. Recent developments further demonstrate that non-invertible symmetries can impose nontrivial constraints on entanglement patterns, influence topological order, and enforce selection rules that transcend those encountered in the setting of ordinary invertible symmetries Casini:2019kex; Copetti:2024dcz; AliAhmad:2025bnd. Motivated by these developments, this paper examines how non-invertible (categorical) symmetries modify entanglement measures in certain two-dimensional topological systems at criticality. For global invertible symmetries, SREE has been extensively studied, allowing the entanglement to be resolved sector by sector. Readers are referred to these reviews Castro-Alvaredo:2024azg. In contrast, symmetry resolution for categorical symmetries Lin:2022dhv; Sierra_2024; Das:2024qdx; Heymann:2024vvf; Choi:2024wfm; Das:2025xyz remains far less explored. In (1+1)(1{+}1)-dimensional CFTs with a categorical symmetry encoded by a fusion category $\mathcal{C}$, symmetry resolution can be formulated in terms of boundary CFT (BCFT) DiGiulio:2022jjd; Kusuki:2023bsp; Northe:2023khz through projections onto categorical sectors. These projections are implemented by symmetric (either strongly or weakly) boundary states on which the corresponding topological defect lines can end. Now, on the replica geometry Calabrese:2009qy, inserting compatible defect endpoints along the branch cut yields well-defined sector probabilities and sector-resolved Rényi and von Neumann entropies. This construction is valid whenever the categorical symmetry is non-anomalous, i.e. free of obstructions to gauging Choi:2023PRD; Seifnashri:2025fgd. Alternatively, there is a twist-field approach that compute directly symmetry-resolved entanglement entropy using charged moments and this method is applied to compute SREE for rational CFT with Lie group symmetry by Calabrese:2021wvi and for certain non-invertible Das:2025xyz.
@@
Interestingly, for 2d CFTs with global internal abelian symmetry, SREE shows universal results in the sense of equipartition of entanglement among all charge sector german3; Castro-Alvaredo:2024azg. On the other hand, in the context of rational CFTs (RCFTs) [^3] It turns out that almost all rational CFTs possess non-invertible symmetry., for non-invertible categorical symmetry-resolved entanglement entropy, this equipartition of the entanglement is broken Kusuki:2023bsp. These deviations are controlled by the fusion data of $\mathcal{C}$ and by the choice of symmetric boundary condition. Note that computation of the entanglement requires partitioning the Hilbert space, hence the boundary state placed at the entangling surface plays a crucial role in refining the entropy Ohmori:2014eia; Cardy_2016. In contrast to the group-like setting, non-invertible symmetries generally admit multiple, and often inequivalent, choices of boundary conditions. Although this issue has been noted in the literature Choi:2024wfm; Choi:2023PRD; Choi:2024tri; Lin:2022dhv; Heymann:2024vvf; Das:2024qdx, previous discussions have largely centered on the so-called weakly/strongly symmetric boundary condition and with a multiplicity-free fusion ring. In this work, we revisit this question through a series of concrete examples that clarify how different boundary conditions naturally arise and how they affect the symmetry resolution. In particular, we highlight situations in which the weakly symmetric boundary state is absent and motivate the use of cloaking boundary conditions (or states) Brehm:2022JPA; Brehm:2024arXiv, which, despite not corresponding to physical Cardy states, play a crucial role in implementing symmetry resolution in non-invertible symmetries consistently.
@@
Most developments on symmetry resolution have focused on unitary settings, where the physical Hilbert space and defect operations admit a positive-definite inner product, with positive central charge. In contrast, symmetry resolution for categorical symmetries in non-unitary setting remains comparatively unexplored, despite clear motivation. By the bulk–boundary correspondence, when a (2+1)(2{+}1)-dimensional topological phase is described by a unitary TQFT, the universal properties of its gapless boundary are captured by a two-dimensional unitary RCFT Witten:1988hf; Moore:1988qv; Moore:1989vd; Dijkgraaf:1989pz; Verlinde:1988sn; Fuchs:2009iz. Extending this logic beyond unitarity is natural Gang:2023rei. It has been argued that a family of non-unitary TQFTs arises from topological twists of certain three-dimensional $\mathcal{N}=4$ rank-zero (S-fold) SCFTs, producing modular data that generalize the non-unitary Haagerup–Izumi modular data Gang:2023ggt; Gang:2022kpe. This, in turn, leads to the question of whether there exist boundary RCFTs compatible with such generalized Haagerup modular data Haagerup1994; Evans:2010yr. The answer is affirmative: using the so-called Riemann–Hilbert approach, conformal characters consistent with the proposed modular data have been explicitly constructed Gang:2023ggt. These developments supply a concrete arena in which to formulate and study categorical symmetry resolved entanglement for non-unitary RCFTs, thereby extending symmetry resolution beyond the unitary paradigm. At this point, note that there are recent studies aimed at constructing critical lattice models in two dimensions, realizing the symmetry associated with the unitary Haagerup fusion category $H_{3}$ Huang:2021nvb; Vanhove:2021zop; Jia:2024wnu; Bottini:2025hri and extracted the central charge (and operator spectra) which turns out to be $c\sim 2$. Then it was further conjectured that this two-dimensional CFT can be constructed by a taking coset of a Haagerup rational CFT characterized central charge multiple of $8$ Evans:2010yr, even though there is a puzzle regarding the existence of such $c=8$ unitary rational CFT with Haagerup fusion $\mathcal{Z}(H_{3})$ (for more details, interested readers are refered to T. Gannon’s talk Gannon:2025birs) [^4] Recently, entanglement asymmetry for $\mathcal{Z}(H_{3})$ has been computed Benini:2025lav.. Nonetheless, we draw motivations from these studies and focus on certain non-unitary generalized Haagerup fusion categories Gang:2023ggt(some of them are tensor products of certain non-unitary minimal models) and discuss symmetry-resolution of entanglement entropy.
@@ 3 Symmetry resolved entanglement entropy (SREE)
where $c_{\rm eff}=c-24\Delta_{\rm GS}$ and $\Delta_{\rm GS}$ is the conformal dimension of the ground state (for unitary theory ground state is the vacuum state with $\Delta_{\rm GS}=0$, but non-zero for non-unitary theory). In this limit, the SRRE for the irreducible block becomes
@@
which applies to a broader class of boundary conditions. Here $\overline{\mathcal{N}}_{\rho B_{1}}^{\,\,B_{2}}$ denotes the boundary fusion coefficient defined via the action of a topological defect line on a boundary state:
@@
For weakly symmetric Cardy boundaries, the action of topological defect lines on boundary Ishibashi states closes according to the bulk fusion algebra:
@@ 5.1 Lee-Yang Minimal Model $\mathcal{M}(5,2)$: a simplest example
where the constant term
@@
$\frac{1}{2}\log\frac{5+\sqrt{5}}{10}=\log\left(\frac{d_{\mathcal{R}_{2}}}{\sum\limits_{\mathcal{R}_{i}}d_{\mathcal{R}_{i}}}\right)+\log(\langle\!\langle\Phi_{2}|\!|\Omega\rangle\!\langle\Omega|\!|\Phi_{2}\rangle\!\rangle).$ (93)
@@ 6 Summary and Conclusion
  • General Framework for Categorical Symmetry Resolution. We first discussed the role of boundary conditions in symmetry resolution and introduced a general prescription, inspired by the SymTFT construction of Choi:2024wfm, for computing symmetry-resolved entanglement entropies in arbitrary categorical charge sectors. This formalism applies not only to weakly and strongly symmetric boundary conditions, but also to the more subtle cloaking boundary conditions, and it naturally accommodates fusion categories with multiplicities. We demonstrated that the resulting formula is universally applicable across a wide range of RCFTs, irrespective of unitarity or diagonality.

### v2

@@ Abstract
We explore several aspects of the categorical symmetry-resolved entanglement entropy (SREE) in two-dimensional Rational Conformal Field Theories (RCFTs) and express it directly in terms of the modular data of the theory. Motivated by Choi:2024wfm, we provide a general formula for SREE that applies to symmetric (weakly/strongly) and cloaking boundary conditions as well as for fusion rings with multiplicities without invoking any SymTFT construction, relying instead on a purely 2d RCFT analysis. We check the formula against several explicit examples. Additionally, we study symmetry resolution for both categorical and invertible symmetries in (non-)diagonal RCFTs and comment on the subtleties that arise in these cases. Finally, we extend our analysis to diagonal non-unitary RCFTs, focusing on theories with generalized Haagerup–Izumi modular data, and find full agreement with the given formula.
@@ 1 Introduction
Symmetries are omnipresent in nature. In physics, symmetry plays a crucial role in understanding the dynamics of quantum field theory and provides a framework for going beyond. According to Wigner Wigner:1959, each symmetry is associated with a unitary operation that has an inverse, i.e., group-like. Gaiotto et al. Gaiotto:2014kfa showed that the generalized global symmetry of a theory is associated with a topological defect operator, and there are plenty of examples found in condensed matter systems that realize this generalized global symmetry McGreevy:2022oyu. One of the classes of these generalized symmetries is famously known as non-invertible symmetry Bhardwaj:2017xup; Chang:2018iay; Thorngren:2019iar; Heidenreich:2021xpr; Thorngren:2021yso; Choi:2022jqy; GarciaEtxebarria:2022vzq; Cordova:2022ieu; Cordova:2022ruw; Choi:2022fgx; Choi:2022rfe; Apruzzi:2022rei; Shao:2023gho; Seiberg:2023cdc; Schafer-Nameki:2023jdn; Bhardwaj:2023wzd; Bhardwaj:2023ayw; Bhardwaj:2023fca; Choi:2023pdp; Lin:2023uvm; Iqbal:2024pee; Choi:2024rjm; Seiberg:2024gek; Bhardwaj:2024wlr; Bhardwaj:2024kvy; Putrov:2024uor; Shao:2025mfj; Shao:2025qvf  [^1] The references mentioned here are by no means exhaustive. Interested readers are referred to some of these reviews and lecture notes and their citations McGreevy:2022oyu; Shao:2023gho; Schafer-Nameki:2023jdn for more references., which can not be described by a unitary operation, more specifically, it does not have an inverse. In mathematics literature, it is known as a category Bartsch:2023wvv; Costa:2024wks. For these kinds of categorical (non-invertible) symmetries, associated topological operators $\{\mathcal{L}_{a}\}$ obey fusion rules of a tensor category rather than a group law, as shown below,
@@
There are numerous examples that demonstrate this kind of categorical symmetry exists in various branches of physics, including high-energy theory, mathematical physics, and condensed matter physics, particularly in low-dimensional quantum systems that are useful for quantum computation Preskill:1999he; Kitaev:2005dm; Kitaev:1997wr; PhysRevLett.96.110405; Aharonov:2005nkd; Feiguin:2006ydp; Dennis:2001nw; McGreevy:2022oyu; PRXQuantum.5.037001; Okada:2024qmk; Li:2024gwx; Fukusumi:2024cnl; Bhardwaj:2024ydc; PhysRevX.15.011058; Li:2025bgo; KNBalasubramanian:2025vum. Before delving into the main discussion of the paper regarding the categorical symmetry resolution of the entanglement entropy, we briefly discuss the representation of categories and the importance of dealing with the Drinfeld centre or the quantum double $Z(\mathcal{C})$ of a given category $\mathcal{C}\,.$
@@
Parallel to recent advances in the characterization of (non-)invertible symmetries, a complementary line of research has focused on the role of symmetries in shaping entanglement structures in quantum field theories (QFTs). Entanglement serves as a fundamental measure of quantum correlations and phases of matter, and provides a particularly sensitive probe of underlying symmetry principles, including generalized and non-invertible symmetries. In this context, symmetry-resolved entanglement entropy (SREE) Goldstein:2017bua; german3; Belin:2013uta; Islam:2015mom; doi:10.1126/science.aau0818; FG; Murciano:2019wdl; Murciano:2020lqq; Parez:2021pgq; Murciano:2020vgh; Zhao:2020qmn; ourPartI; Horvath:2021rjd; Weisenberger:2021eby; Magan:2021myk; Ares:2022gjb; DiGiulio:2022jjd; Zhao:2022wnp; Gaur:2023yru; Kusuki:2023bsp; Northe:2023khz; Gaur:2024vdh; Banerjee:2024ldl [^2] For more references and details, interested readers are referred to the following review of symmetry-resolved entanglement measure Castro-Alvaredo:2024azg. has emerged as a refined diagnostic that decomposes the total entanglement into contributions associated with individual symmetry sectors, thereby elucidating how symmetry structures govern the distribution of quantum information. Recent developments further demonstrate that non-invertible symmetries can impose nontrivial constraints on entanglement patterns, influence topological order, and enforce selection rules that transcend those encountered in the setting of ordinary invertible symmetries Casini:2019kex; Lin:2022dhv; Copetti:2024dcz; AliAhmad:2025bnd. Motivated by these developments, this paper examines how non-invertible (categorical) symmetries modify entanglement measures in certain two-dimensional topological systems at criticality. For global invertible symmetries, SREE has been extensively studied, allowing the entanglement to be resolved sector by sector. Readers are referred to these reviews Castro-Alvaredo:2024azg. In contrast, symmetry resolution for categorical symmetries Sierra_2024; Benedetti:2024dku; Das:2024qdx; Heymann:2024vvf; Choi:2024wfm; Das:2025xyz remains far less explored. In (1+1)(1{+}1)-dimensional CFTs with a categorical symmetry encoded by a fusion category $\mathcal{C}$, symmetry resolution can be formulated in terms of boundary CFT (BCFT) DiGiulio:2022jjd; Kusuki:2023bsp; Northe:2023khz through projections onto categorical sectors. These projections are implemented by symmetric (either strongly or weakly) boundary states on which the corresponding topological defect lines can end. Now, on the replica geometry Calabrese:2009qy, inserting compatible defect endpoints along the branch cut yields well-defined sector probabilities and sector-resolved Rényi and von Neumann entropies. This construction is valid whenever the categorical symmetry is non-anomalous, i.e. free of obstructions to gauging Choi:2023PRD; Seifnashri:2025fgd. Alternatively, there is a twist-field approach that compute directly symmetry-resolved entanglement entropy using charged moments and this method is applied to compute SREE for rational CFT with Lie group symmetry by Calabrese:2021wvi and for certain non-invertible Das:2025xyz.
@@
Interestingly, for 2d CFTs with global internal abelian symmetry, SREE shows universal results in the sense of equipartition of entanglement among all charge sector german3; Castro-Alvaredo:2024azg  [^3] One may note that for abelian symmetry the entanglement equipartition does not hold for all order of UV cutoff, see Northe:2023khz; Northe:2025qcv.. On the other hand, in the context of rational CFTs (RCFTs) [^4] It turns out that almost all rational CFTs possess non-invertible symmetry., for non-invertible categorical symmetry-resolved entanglement entropy, this equipartition of the entanglement is broken Kusuki:2023bsp. These deviations are controlled by the fusion data of $\mathcal{C}$ and by the choice of symmetric boundary condition. Note that computation of the entanglement requires partitioning the Hilbert space, hence the boundary state placed at the entangling surface plays a crucial role in refining the entropy Ohmori:2014eia; Cardy_2016. In contrast to the group-like setting, non-invertible symmetries generally admit multiple, and often inequivalent, choices of boundary conditions. Although this issue has been noted in the literature Choi:2024wfm; Choi:2023PRD; Choi:2024tri; Lin:2022dhv; Heymann:2024vvf; Das:2024qdx, previous discussions have largely centered on the so-called weakly/strongly symmetric boundary condition and with a multiplicity-free fusion ring. In this work, we revisit this question through a series of concrete examples that clarify how different boundary conditions naturally arise and how they affect the symmetry resolution. In particular, we highlight situations in which the weakly symmetric boundary state is absent and motivate the use of cloaking boundary conditions (or states) Brehm:2022JPA; Brehm:2024arXiv, which, despite not corresponding to physical Cardy states, play a crucial role in implementing symmetry resolution in non-invertible symmetries consistently.
@@
Most developments on symmetry resolution have focused on unitary settings, where the physical Hilbert space and defect operations admit a positive-definite inner product, with positive central charge. In contrast, symmetry resolution for categorical symmetries in non-unitary setting remains comparatively unexplored, despite clear motivation Fossati:2023zyz; Fukusumi:2025xrj. By the bulk–boundary correspondence, when a (2+1)(2{+}1)-dimensional topological phase is described by a unitary TQFT, the universal properties of its gapless boundary are captured by a two-dimensional unitary RCFT Witten:1988hf; Moore:1988qv; Moore:1989vd; Dijkgraaf:1989pz; Verlinde:1988sn; Fuchs:2009iz. Extending this logic beyond unitarity is natural Gang:2023rei. It has been argued that a family of non-unitary TQFTs arises from topological twists of certain three-dimensional $\mathcal{N}=4$ rank-zero (S-fold) SCFTs, producing modular data that generalize the non-unitary Haagerup–Izumi modular data Gang:2023ggt; Gang:2022kpe. This, in turn, leads to the question of whether there exist boundary RCFTs compatible with such generalized Haagerup modular data Haagerup1994; Evans:2010yr. The answer is affirmative: using the so-called Riemann–Hilbert approach, conformal characters consistent with the proposed modular data have been explicitly constructed Gang:2023ggt. These developments supply a concrete arena in which to formulate and study categorical symmetry resolved entanglement for non-unitary RCFTs, thereby extending symmetry resolution beyond the unitary paradigm. At this point, note that there are recent studies aimed at constructing critical lattice models in two dimensions, realizing the symmetry associated with the unitary Haagerup fusion category $H_{3}$ Huang:2021nvb; Vanhove:2021zop; Jia:2024wnu; Bottini:2025hri and extracted the central charge (and operator spectra) which turns out to be $c\sim 2$. Then it was further conjectured that this two-dimensional CFT can be constructed by a taking coset of a Haagerup rational CFT characterized central charge multiple of $8$ Evans:2010yr, even though there is a puzzle regarding the existence of such $c=8$ unitary rational CFT with Haagerup fusion $\mathcal{Z}(H_{3})$ (for more details, interested readers are refered to T. Gannon’s talk Gannon:2025birs) [^5] Recently, entanglement asymmetry for $\mathcal{Z}(H_{3})$ has been computed Benini:2025lav.. Nonetheless, we draw motivations from these studies and focus on certain non-unitary generalized Haagerup fusion categories Gang:2023ggt(some of them are tensor products of certain non-unitary minimal models) and discuss symmetry-resolution of entanglement entropy.
@@ 3 Symmetry resolved entanglement entropy (SREE)
where $c_{\rm eff}=c-24\Delta_{\rm GS}$ and $\Delta_{\rm GS}$ is the conformal dimension of the ground state (for unitary theory ground state is the vacuum state with $\Delta_{\rm GS}=0$, but non-zero for non-unitary theory). In this limit, the SREE for the irreducible block becomes
@@
which applies to a broader class of boundary conditions. Here $\overline{\mathcal{N}}_{\rho B_{1}}^{\,\,B_{2}}$ denotes the boundary fusion coefficient defined via the action of a topological defect line on a boundary state [^10] We propose this interpretation of the working formula without invoking the SymTFT construction, relying instead on a purely 2d RCFT analysis, in contrast to Choi:2024wfm. :
@@
For weakly symmetric Cardy boundaries, the action of topological defect lines on boundary Ishibashi states closes according to the bulk fusion algebra Graham:2003nc  [^11] Interested readers are referred to Fukusumi:2020irh for some applications of these boundary conditions in the context of Ising model for describing edge modes.:
@@ 5.1 Lee-Yang Minimal Model $\mathcal{M}(5,2)$: a simplest example
where the constant term  [^18] For non-unitary theory, bra and ket are not complex conjugate of each other in general. But, in our case it does and our results matches with those in Fukusumi:2025xrj. We thank Yoshiki Fukusumi for pointing out thid connection to us.
@@
$\frac{1}{2}\log\frac{5+\sqrt{5}}{10}=\log\left(\frac{d_{\mathcal{R}_{2}}}{\sum\limits_{\mathcal{R}_{i}}d_{\mathcal{R}_{i}}}\right)+\log(\langle\!\langle\Phi_{2}|\!|\Omega\rangle\!\langle\Omega|\!|\Phi_{2}\rangle\!\rangle)$ (93)
@@ 6 Summary and Conclusion
  • General Framework for Categorical Symmetry Resolution. We first discussed the role of boundary conditions in symmetry resolution and introduced a general prescription of computing SREE. Inspired by Choi:2024wfm, we provide a general formula for the SREE that uniformly applies to symmetric (weakly or strongly) and cloaking boundary conditions, as well as to fusion rings with multiplicities, without invoking any SymTFT construction. Our approach relies solely on a purely two-dimensional RCFT analysis. This formalism applies not only to weakly and strongly symmetric boundary conditions, but also to the more subtle cloaking boundary conditions, and it naturally accommodates fusion categories with multiplicities. We demonstrated that the resulting formula is universally applicable across a wide range of RCFTs, irrespective of unitarity or diagonality.